Institutions ... but which?

The importance of the institutional setting in determining the trade between EU member states.

Joanna Tyrowicz

May 10^{th}, 2004

Abstract

Does trade stimulate GDP or is it the other way around? Not only is the answer to this question crucial for policy implications, but it also helps to understand the nature of wealth creation. Number of studies confirmed that overall openness stimulates economic growth allowing also to explain differences in income levels (cfr. the works of Xavier Sala-i-Martin, Jeffrey Frankel and Andrew K. Rose). However, although for obvious reasons trade and GDP seem closely related, one was so far unable to unambiguously ascertain, what the causality is. Nevertheless, it should not discourage economists from analysing what stands behind the international exchange of goods and services.

Analysing the trade patterns between European Union member states seems particularly important for at least few reasons. Primarily, EU is the most integrated regional trade block in the world giving valuable experience of what such integration really means. Observing which factors determine patterns of trade seems therefore interesting.

Secondly, the extent to which EU member states trade with each other affects the level of their integration. Thus, since the asymmetry of external shock response might be marginalised by the deeper integration, it might be crucial for reducing the risk of EU collapse. For that reason, the determinants of trade are also important for other integration levels, as common market and political union.

In addition, the Intra-EU trade patterns determinants are particularly important for Poland and the other accessing countries. The process of adjusting the institutions to the EU standards, originated several years ago shall not be ended on the accession day. Thus, understanding which institutions are important for stimulating the Intra-EU exchange remains an important issue.

In this paper we aim at identifying the institutional determinants of trade. It has been argued frequently, that institutions, geography are together important stimuli for economic growth. Nevertheless, as shown by Dani Rodrik, some institutions can influence GDP levels indirectly and in that respect, trade was found to be an important transmission channel. Unlike other studies, however, we intend to analyse which institutions exactly can be considered essential.

The main goal of this paper is to identify and analyse main institutional determinants of the intra-EU trade. We try to answer the following question: which institutions help two companies to trade, despite their location? In order to achieve this, a gravity equation was constructed where formal and informal institutions were explicitly specified. The reached conclusions point to the importance of the legal framework, its quality as well as the prevailing and intense home bias effect.